# Hybrid Quantum-inspired Resnet and Densenet for Pattern Recognition

### v6

@@ Abstract
In this paper, we propose two hybrid quantum-inspired neural networks with adaptive residual and dense connections respectively for pattern recognition. We explain the frameworks of the symmetrical circuit models in the quantum-inspired layers in our hybrid models. We also illustrate the potential superiority of our hybrid models to prevent gradient explosion owing to the quantum-inspired layers. Groups of numerical experiments on generalization power show that our hybrid models possess roughly the same level of generalization power as the pure classical models with different noisy datasets utilized. Furthermore, the comparison on generalization ability between our hybrid models and a state-of-the-art hybrid quantum-classical convolutional network demonstrates 3%-4% higher accuracy of our hybrid densely-connected model than the hybrid quantum-classical network. Simultaneously, in terms of groups of experiment on robustness, the results demonstrate that our two hybrid models outperform pure classical models notably in resistance to parameter attacks with various asymmetric noises. They also indicate the superiority of our densely-connected hybrid model over the hybrid quantum-classical network under both symmetrical and asymmetrical attacks. Furthermore, an ablation study indicate that the recognition accuracy of our two hybrid models is 2%-3% higher than that of the traditional quantum-inspired neural network without residual or dense connection. Eventually, we discuss the application scenarios of our hybrid models by analyzing their computational complexities.
@@ 1 Introduction
As the cornerstone of AI technology, deep neural network algorithms have played a pivotal role over the past few decades [1, 2, 3, 4, 5, 6, 7]. Typical models incorporate deep residual and dense networks (Resnet and Densenet) in image recognition, Large Language Models (LLM) in natural language processing, and Sora model in multi-modal generations [8, 9, 10, 11]. Nevertheless, due to the limited generalization power and robustness of classical neural networks, the design of neural networks with stronger generalization power and robustness over pure classical models is tremendously expected nowadays [12, 13, 14, 15, 16]. Simultaneously, quantum-inspired neural networks, which obey the laws of quantum computing and deep learning, appear with their prominent performances on certain circumstances [17, 18, 19]. Innovative frameworks include quantum-inspired stochastic walks and variational circuit models, etc [19, 20, 21, 22]. What’s more, to improve the universality of the quantum-inspired models under complex environments, hybrid quantum-inspired neural networks have been explored in some studies [23, 24, 25, 26]. Connected with classical convolutional layers, they demonstrate remarkable performance on image detection and other tasks[27, 28, 29]. Comprised of the pure classical part and the quantum-inspired part, they could leverage different advantages from both pure classical and quantum-inspired neural networks [25, 26].
Nevertheless, despite a few works of hybrid quantum-inspired neural networks, there is a lack of them which leverage the amalgamation of circuit models and residual or dense connection [8, 9, 30]. Hence, we firstly design the hybrid Quantum-inspired Residual Feedforward Neural Network (QRFNN) and hybrid Quantum-inspired Dense Feedforward Neural Network (QDFNN). We also devise hybrid Quantum-inspired Residual Convolutional Neural Networks (QRCNN) and hybrid Quantum-inspired Dense Convolutional Neural Networks (QDCNN) for image problems. We apply our four hybrid models on pattern classification problems. A Hybrid Quantum-classical convolutional Network (HQNet), classical Multi-Layer Perceptrons (MLPs) and Convolutional Neural Networks (CNNs) with detailed structures serve as state-of-the-arts for model comparison. Hence, in the paper, we:
@@
    conduct assorted numerical experiments to compare our hybrid models with a state-of-the-art hybrid quantum-classical convolutional network and pure classical models;
@@
Next, in section 2, we review the corresponding works of classical residual and dense frameworks and quantum-inspired neural networks with circuit models. In section 3, we introduce some basic concepts of circuit models, residual structure and dense structure [8, 9, 30]. As for section 4, we explain our hybrid models including the layer details and parameter learning details. Following it, some numerical experiments are implemented in section 5, which involve comparisons between our hybrid models and pure classical and the hybrid quantum-classical models. We also do an ablation study incorporating comparisons between our hybrid models and the quantum-inspired neural network without residual or dense connection in section 5. Finally, we illustrate the application scenarios and some future work of our hybrid models in section 6.
@@ 2.2 Quantum-inspired neural networks with circuit models
Quantum-inspired neural networks, a classical algorithm essentially, are impacted by some quantum laws [21, 22]. A typical model is circuit model [22, 42]. In the authentic quantum circuit system, classical data is encoded into a quantum state. And the data feature is acquired via evolutionary processes and stored in the parameters in the unitary gates of the circuit models. And the design of the quantum-inspired part of our hybrid models draw on the frameworks of the circuit models in authentic quantum system [22]. Moreover, our quantum-inspired layers only follows some laws of circuit models, therefore, our hybrid models require more improvements to be deployed on authentic quantum computers. However, in previous studies of circuit models, the architectures of the circuits are usually unsymmetrical, which may result in information loss [22]. Moreover, the unitary gates utilized incorporate ${\rm R}_{\rm Y},{\rm R}_{\rm Z}$, Hadamard and CNOT gates, some of which contain no parameter and could not assist feature learning [25, 26, 42]. Hence, to enable our models to fully extract the feature, each gate in the quantum-inspired part contains a parameter. Moreover, one novelty lies in the proposal of the symmetrical ”V” shape in each quantum-inspired layer of our hybrid models in Fig.3, which may also help feature storage [19, 43]. In addition, traditional quantum-inspired networks may not possess very outstanding learning capability [21], thus another novelty lies in the combination of the adaptive residual and dense connections and the symmetrical circuit models.
@@ 3.1 Circuit models
where $U_{m}$ represents the $(m+1)^{th}$ unitary gate in the $m^{th}$ evolution. And |$\rm{\ket{\psi_{0}}}$ means the initial state. The gate utilized in the paper, $T_{i}(\theta^{l-1}_{r})$:
={pNiceMatrix}$\scriptsize=\pNiceMatrix[first-col,last-row]&10\cdots\cdots\cdots\cdots\cdots 0\\ 01\vdots\\ \vdots\ddots\vdots\\ i\vdots\cos({\mathrm{\theta}}^{l-1}_{r}){\rm-sin}(\theta^{l-1}_{r})\vdots\\ i+1\vdots\sin({\rm\theta}^{l-1}_{r})\cos({\rm\theta}^{l-1}_{r})\vdots\\ \vdots\ddots\vdots\\ \vdots 10\\ 0\cdots\cdots\cdots\cdots\cdots 01\\ ii+1\$ (3)
As for Eq.(3), $i$ denotes the row index of the term ${\rm-sin}({\theta}^{l-1}_{r})$, also the column index of the term $\sin({\theta}^{l-1}_{r})$ $(i\in\mathbb{N}^{+})$. And $\theta^{l-1}_{r}$ means the $(r+1)^{th}$ parameter in the $l^{th}$ layer $(r\in\mathbb{N},l\in\mathbb{N}^{+})$. And $r$ has relations with $i$, which is shown in Eq.(4).
@@ 3.2 Resnet and Densenet learning
Figure 3: QRFNN and QDFNN. (a) QRFNN. (b) QDFNN.
@@ 4.1 Layer details of QRFNN and QDFNN
Fig.3 demonstrates the QRFNN and QDFNN structures of 2 qubits with 2 layers, where each line represents one-dimensional vector space of a quantum state. The green, blue circles denote the input and output neurons respectively. To consider the general occasions, suppose we build $L$ quantum-inspired layers, the quantum-inspired part, as hidden layers, in QRFNN and QDFNN rather than MLP hidden layers with neurons. So there are $L+2$ layers totally in both QRFNN and QDFNN. We set the quantum states in the quantum-inspired layers are $N$-dimensional with $n$ qubits utilized for QRFNN and QDFNN $(2^{n-1}\leq N\leq 2^{n})$. And $T_{i}(\theta^{l}_{r})\in\mathbb{R}^{N\times N}$. We also suppose the dimension of the input, $X=(x_{0},x_{1},...,x_{N-1})^{\top}$, is $N$. In terms of $r$ and $i$ in Eq.(3) in this section, we also have:
$r=\begin{cases}i-1&\text{ $0\leq r\leq N-2$ }\\ 2N-3-i&\text{ $N-2<r\leq 2N-4$ }\end{cases}$ (4)
@@
$\prod\limits_{\begin{subarray}{c}j=0\end{subarray}}^{\begin{subarray}{c}j=N\end{subarray}}T_{j+1}(\alpha_{j}^{l-1})\coloneqq T_{1}(\alpha_{0}^{l-1})T_{2}(\alpha_{1}^{l-1})...T_{N+1}(\alpha_{N}^{l-1})$ (5a)
$\prod\limits_{\begin{subarray}{c}j=0\end{subarray}}^{\begin{subarray}{c}j=N\end{subarray}}T_{j+1}(\beta_{j}^{l-1})\coloneqq T_{1}(\beta_{0}^{l-1})T_{2}(\beta_{1}^{l-1})...T_{N+1}(\beta_{N}^{l-1})$ (5b)
@@
Figure 4: General structure of the $l^{th}$ quantum-inspired layer. Here the parameter is represented as $\theta$, but in QRFNN and QDFNN, we use $\alpha$ and $\beta$ as parameters respectively. And each line represents one-dimensional vector space of the states. Furthermore, the ”V” shape helps feature learning of each category of the dataset. On one hand, if the circuit shape is not the symmetrical ”V” shape, there could be imbalance of the number of parameters of each category of a dataset for feature learning. On the other hand, each unitary gate is a sparse matrix with only one parameter, which ensure the convienience of computation and guarantee that over-parameterization does not occur.
@@
and |$\ket{\psi_{l}},\ket{\psi_{l}}^{\prime}$ are respectively the output states of the $l^{th}$ quantum-inspired layers of QRFNN and QDFNN. $\oplus$ denotes the element-wise addition mechanism. The terms $U_{l-1}\ket{\psi_{l-1}}$ and $U_{l-1}^{\prime}\ket{\psi_{l-1}}^{\prime}$ indicate the quantum-inspired parts of the two models obey laws of circuit models [22]. However, because our hybrid models also follow the deep learning laws, |$\ket{\psi_{l}}$ and |$\ket{\psi_{l}}^{\prime}$ can be also respectively represented as column vectors $O_{l}$ and $O_{l}^{\prime}$ (see supplementary materials part 2). For QRFNN, $O_{l}=(o_{l,0},o_{l,1},...,o_{l,N-1})^{\top}$. For QDFNN, $O_{l}^{\prime}=(o_{l,0}^{\prime},o_{l,1}^{\prime},...,o_{l,N-1}^{\prime})^{\top}$. And $o_{l,j},o_{l,j}^{\prime}$ denote the output at the $(j+1)^{th}$ dimensional index of the $l^{th}$ quantum-inspired hidden layer of the two models respectively. And we have:
@@
where $\sigma$ denotes any activation function, $W_{L+1}$ and $W_{L+1}^{\prime}$ mean the weight matrices of the output layers of QRFNN and QDFNN separately, while $B_{L+1},B_{L+1}^{\prime}$ represent biases of QRFNN and QDFNN respectively. Therefore, our hybrid neural networks are comprised of quantum-inspired hidden layers, as well as pure classical input and output layers.

### v7

@@ Abstract
In this paper, we propose two hybrid quantum-inspired neural networks with adaptive residual and dense connections respectively for pattern recognition. We explain the frameworks of the symmetrical circuit models in the quantum-inspired layers in our hybrid models. We also illustrate the potential superiority of our hybrid models to prevent gradient explosion owing to the sine and cosine functions in the quantum-inspired layers. Groups of numerical experiments on generalization power showcase that our hybrid models are comparable to the pure classical models with different noisy datasets utilized. Furthermore, the comparison between our hybrid models and a state-of-the-art hybrid quantum-classical convolutional network demonstrates 3%-4% higher accuracy of our hybrid densely-connected model than the hybrid quantum-classical network. Additionally, compared with other two hybrid quantum-inspired residual networks, our hybrid models showcase a little higher accuracy on image datasets with asymmetrical noises. Simultaneously, in terms of groups of robustness experiments, the outcomes demonstrate that our two hybrid models outperform pure classical models notably in resistance to adversarial parameter attacks with various asymmetrical noises. They also indicate the slight superiority of our densely-connected hybrid model over the hybrid quantum-classical network to both symmetrical and asymmetrical attacks. Meanwhile, the accuracy of our two hybrid models is a little bit higher than that of the two hybrid quantum-inspired residual networks. In addition, an ablation study indicate that the recognition accuracy of our two hybrid models is 2%-3% higher than that of the traditional quantum-inspired neural network without residual or dense connection. Eventually, we discuss the application scenarios of our hybrid models by analyzing their computational complexity.
@@ 1 Introduction
As the cornerstone of AI technology, deep neural network algorithms have played a pivotal role over the past few decades [1, 2, 3, 4, 5, 6, 7]. Typical models incorporate deep residual and dense networks (Resnet and Densenet) in image recognition, Large Language Models (LLM) in natural language processing, and Sora model in multi-modal generations [8, 9, 10, 11]. Nevertheless, due to the limited generalization power and robustness of classical neural networks, novel neural networks with stronger generalization power under assorted environments and robustness to adversarial attacks over pure classical models is tremendously expected nowadays [12, 13, 14, 15, 16]. Simultaneously, quantum-inspired neural networks, which obey the laws of quantum computing and deep learning, appear with their prominent performances on certain circumstances [17, 18, 19]. Innovative frameworks incorporate quantum-inspired stochastic walks and variational circuit models, etc [19, 20, 21, 22]. While stochastic walks incorporate the solution and evolution of Hamiltonian, variational circuit models involve continuous multiplication of unitary matrices. What’s more, to integrate the upsides of pure classical neural networks with that of quantum-inspired networks, hybrid quantum-inspired neural networks have been explored in some studies [23, 24, 25, 26]. Amalgamated with classical convolutional layers, they demonstrate remarkable performance on image detection and other tasks [27, 28, 29]. Comprised of the pure classical part and the quantum-inspired part, they are applicable to more types of problems [25, 26].
Nevertheless, despite a few works of hybrid quantum-inspired models, the experimental results of them showcase that those hybrid quantum-inspired networks only either show higher accuracy in pattern classification problems, or possess stronger generalization ability in different scenarios, or are more robust than traditional classical or quantum neural networks. Additionally, models with distinguished robustness possess lower accuracy or weaker generalization ability than classical ones [23, 25, 26, 27, 28, 29]. There has been no hybrid quantum-inspired models possessing both high accuracy and enhanced generalization capability as well as robustness to adversarial attacks. Meanwhile, while several quantum neural networks with circuit models demonstrate unique robustness to noise attacks, numerous classical residual and dense neural networks showcase exceptional accuracy and generalization power [2, 3, 8, 9, 21, 22]. What’s more, a common problem on quantum neural networks is barren plateau, which could be avoided by the addition of residual frameworks [23]. Concurrently, there is a lack of hybrid quantum-inspired models leveraging the amalgamation of circuit models and residual or dense connection [8, 9, 30]. Hence, to facilitate the universality and comprehensiveness of hybrid quantum-inspired models under complex environments, we firstly devise the hybrid Quantum-inspired Residual Feedforward Neural Network (QRFNN) and hybrid Quantum-inspired Dense Feedforward Neural Network (QDFNN). We also devise hybrid Quantum-inspired Residual Convolutional Neural Networks (QRCNN) and hybrid Quantum-inspired Dense Convolutional Neural Networks (QDCNN) for image problems. We apply our four hybrid models on pattern classification problems. A Hybrid Quantum-classical convolutional Network (HQNet), pure classical Multi-Layer Perceptrons (MLPs), Convolutional Neural Networks (CNNs) and two hybrid Quantum-inspired Residual neural Networks (ResQNet1 and ResQNet2) with detailed structures serve as state-of-the-arts for model comparison. Hence, in the paper, we:
@@
    conduct assorted numerical experiments to compare our hybrid models with a state-of-the-art hybrid quantum-classical convolutional network, pure classical models and two hybrid quantum-inspired residual networks;
@@
The contributions of this paper are that:
  • 1.
    we devise hybrid quantum-inspired neural networks with proposed adaptive residual or dense connection. We also design symmetrical ”V” circuits with $N-2$ unitary sparse matrices forming one quantum-inspired layer as the residual or dense block, where $N$ is the dimension of the Hilbert space in the quantum-inspired layers;
  • 2.
    our densely-connected hybrid models marginally outperform the state-of-the-art HQNet on generalization power when the datasets contain symmetrical or asymmetrical noises;
  • 3.
    our two hybrid models demonstrate a little bit higher accuracy than other two state-of-the-art hybrid quantum-inspired residual networks to asymmetrical noise attacks or on image classification when the datasets contain asymmetrical noises;
  • 4.
    our hybrid models surpass classical MLPs and CNNs a lot on robustness to asymmetrical noise attacks;
  • 5.
    by analyzing the loss functions and parameter gradients of classical and our hybrid models, the reason why our hybrid models could resist asymmetrical noise attacks lies in their capability to prevent gradient explosion.
Next, in section 2, we review the corresponding works of classical residual and dense frameworks and quantum-inspired neural networks with circuit models. In section 3, we introduce some basic concepts of circuit models, residual and dense structures [8, 9, 30]. As for section 4, we explain our hybrid models including the layer details and parameter learning details. Following it, numerous experiments are implemented in section 5, which involve comparisons between our hybrid models and pure classical models, the hybrid quantum-classical models and other two hybrid quantum-inspired residual networks. We also conduct an ablation study incorporating comparisons between our hybrid models and the quantum-inspired neural network without residual or dense connection in section 5. Finally, we illustrate the application scenarios and some future works of our hybrid models in section 6.
@@ 2.2 Quantum-inspired neural networks with circuit models
Quantum-inspired neural networks, a classical algorithm essentially, are impacted by some quantum laws [21, 22]. A typical model is circuit model [22, 42]. In the authentic quantum circuit system, classical data is encoded into a quantum state. And the data feature is acquired via evolutionary processes and stored in the parameters in the unitary gates of the circuit models. And the design of the quantum-inspired part of our hybrid models draw on the frameworks of the circuit models in authentic quantum system [22]. Moreover, our quantum-inspired layers only follows some laws of circuit models, therefore, our hybrid models require more improvements to be deployed on authentic quantum computers. However, in previous studies of circuit models, the architectures of the circuits are usually asymmetrical, which may result in information loss [22]. Moreover, the unitary gates utilized incorporate ${\rm R}_{\rm Y},{\rm R}_{\rm Z}$, Hadamard and CNOT gates, some of which contain no parameter and could not assist feature learning [25, 26, 42]. Hence, to enable our models to fully extract the feature, each gate in the quantum-inspired part contains a parameter. Moreover, one novelty lies in the proposal of the symmetrical ”V” shape in each quantum-inspired layer of our hybrid models in Fig.3, which may also help feature storage [19, 43]. In ref. [44], the authors compare the accuracy of symmetrical and asymmetrical circuits for data recognition. For each category of the data, symmetrical circuits possess enough parameters to learn their features. In contrast, asymmetrical circuits may cause imbalance in feature learning, which means they are more efficient and accurate in extracting features from certain categories of a dataset than from other categories of the same dataset[2, 44]. In ref. [43], experiments have showcased prominent accuracy of the ”V” shape network combined with residual connections in image classification problems as well. In addition, traditional quantum-inspired networks may not possess very outstanding learning capability [21], thus another novelty lies in the combination of the adaptive residual and dense connections and the symmetrical circuit models.
@@ 3.1 Circuit models
where $U_{m}$ represents the $(m+1)^{th}$ unitary gate in the $m^{th}$ evolution. And |$\rm{\ket{\psi_{0}}}$ means the initial state.The gate $T_{i}(\theta^{l-1}_{r})$, frequently used in the paper, can be written in the computational basis as the following matrix
{pNiceMatrix}$\scriptsize\pNiceMatrix[first-col,last-row]&10\cdots\cdots\cdots\cdots\cdots 0\\ 01\vdots\\ \vdots\ddots\vdots\\ i\vdots\cos({\mathrm{\theta}}^{l-1}_{r}){\rm-sin}(\theta^{l-1}_{r})\vdots\\ i+1\vdots\sin({\rm\theta}^{l-1}_{r})\cos({\rm\theta}^{l-1}_{r})\vdots\\ \vdots\ddots\vdots\\ \vdots 10\\ 0\cdots\cdots\cdots\cdots\cdots 01\\ ii+1\\ .$ (3)
As for Eq.(3), $i$ denotes the row index of the term ${\rm-sin}({\theta}^{l-1}_{r})$, also the column index of the term $\sin({\theta}^{l-1}_{r})$ $(i\in\mathbb{N}^{+})$. And $\theta^{l-1}_{r}$ means the $(r+1)^{th}$ parameter in the $l^{th}$ layer $(r\in\mathbb{N},l\in\mathbb{N}^{+})$. And $r$ has relations with $i$, shown in Eq.(4).
@@ 3.2 Resnet and Densenet learning
Figure 3: Frameworks of QRFNN and QDFNN. (a) QRFNN. (b) QDFNN.
@@ 4.1 Layer details of QRFNN and QDFNN
Fig.3 demonstrates the QRFNN and QDFNN structures of 2 qubits with 2 layers, where each line represents one-dimensional vector space of a quantum state. The green, blue circles denote the input and output neurons respectively. To contemplate the general occasions, suppose we build $L$ quantum-inspired layers, the quantum-inspired part, as hidden layers, in QRFNN and QDFNN rather than MLP hidden layers with neurons. So there are $L+2$ layers totally in both QRFNN and QDFNN. We set the quantum states in the quantum-inspired layers are $N$-dimensional with $n$ qubits utilized for QRFNN and QDFNN $(2^{n-1}\leq N\leq 2^{n})$. And $T_{i}(\theta^{l}_{r})\in\mathbb{R}^{N\times N}$. We also suppose the dimension of the input, $X=(x_{0},x_{1},...,x_{N-1})^{\top}$, is $N$. In terms of $r$ and $i$ in Eq.(3) in this section, we also have:
$r=\begin{cases}i-1&\text{ $0\leq r\leq N-2$ }\\ 2N-3-i&\text{ $N-2<r\leq 2N-4.$ }\end{cases}$ (4)
@@
$\prod\limits_{\begin{subarray}{c}j=0\end{subarray}}^{\begin{subarray}{c}j=N\end{subarray}}T_{j+1}(\alpha_{j}^{l-1})\coloneqq T_{1}(\alpha_{0}^{l-1})T_{2}(\alpha_{1}^{l-1})...T_{N+1}(\alpha_{N}^{l-1}),$ (5a)
$\prod\limits_{\begin{subarray}{c}j=0\end{subarray}}^{\begin{subarray}{c}j=N\end{subarray}}T_{j+1}(\beta_{j}^{l-1})\coloneqq T_{1}(\beta_{0}^{l-1})T_{2}(\beta_{1}^{l-1})...T_{N+1}(\beta_{N}^{l-1}).$ (5b)
@@
Figure 4: General structure of the $l^{th}$ quantum-inspired layer. Here the parameter is represented as $\theta$, but in QRFNN and QDFNN, we use $\alpha$ and $\beta$ as parameters respectively. And each line represents one-dimensional vector space of the states. Furthermore, the ”V” shape helps feature learning of each category of the dataset. On one hand, if the circuit shape is not the symmetrical ”V” shape, there could be imbalance of the number of parameters of each category of a dataset for feature learning. On the other hand, each unitary gate is a sparse matrix with only one parameter, which ensure the convenience of computation and guarantee that over-parameterization does not occur.
@@
and |$\ket{\psi_{l}},\ket{\psi_{l}}^{\prime}$ are respectively the output states of the $l^{th}$ quantum-inspired layers of QRFNN and QDFNN. $\oplus$ denotes the element-wise addition mechanism. The terms $U_{l-1}\ket{\psi_{l-1}}$ and $U_{l-1}^{\prime}\ket{\psi_{l-1}}^{\prime}$ indicate the quantum-inspired parts of the two models obey laws of circuit models [22]. However, because our hybrid models also follow the deep learning laws, |$\ket{\psi_{l}}$ and |$\ket{\psi_{l}}^{\prime}$ can be also respectively represented as column vectors $O_{l}$ and $O_{l}^{\prime}$ (see supplementary materials section 2). For QRFNN, $O_{l}=(o_{l,0},o_{l,1},...,o_{l,N-1})^{\top}$. For QDFNN, $O_{l}^{\prime}=(o_{l,0}^{\prime},o_{l,1}^{\prime},...,o_{l,N-1}^{\prime})^{\top}$. And $o_{l,j},o_{l,j}^{\prime}$ denote the output at the $(j+1)^{th}$ dimensional index of the $l^{th}$ quantum-inspired hidden layer of the two models respectively. And we have:
@@
where $\sigma$ denotes any activation function, $W_{L+1}$ and $W_{L+1}^{\prime}$ mean the weight matrices of the output layers of QRFNN and QDFNN separately, while $B_{L+1},B_{L+1}^{\prime}$ represent biases of QRFNN and QDFNN respectively. Therefore, our hybrid neural networks are constituted by quantum-inspired hidden layers, as well as pure classical input and output layers.
Besides that, MNIST, FASHIONMNIST and CIFAR image datasets are utilized for validation of QRCNN and QDCNN. Compared with QRFNN and QDFNN, they have one more convolutional module, a few convolutional layers with parallel spatial and channel attentions respectively. Hence, high dimensional features of the image datasets are firstly extracted by parameters in convolutional kernels and pooling operation, which reduces the feature dimension of the data. Then the tensor data is flattened to vectors before they are transferred to quantum-inspired layers. The vectors are encoded into the amplitudes of the quantum states. After the feature learning of quantum-inspired layers, the vectors are transmitted to the output layers. Section 6 of the supplementary materials provides detailed frameworks of QRCNN and QDCNN. The discrepancy between the CNNs and QRCNN only lies in the fully-connected layers in CNNs and the quantum-inspired layers in QRCNN.